OpenStudy (anonymous):

simplify this equation for a line y-6=-4(x-3) give your answer in the form y =_x+_

OpenStudy (anonymous):

\[y-6=-4(x-3)\] remove parentheses using distributive law \[y-6=-4x+12\] add 6 to both sides \[y=-4x+18\]